write lewis structures for these molecules or ions. show formal charge for every atom in the structure. if there are resonance draw them CH3CL, SO3^2-, NH4^+, PO4^3-, CIO3^_, CH3CN
Answer #1

Bamding one Pair o eteuhro n Foumal Valun eumn СHЗC1 Cental arom HC: C: No Resonanu in CHCI thraunge em C = eaun atom Fil en0- No Rson anu NHY HN.H 5 - 0-1183 F.L on Hy du gs F-C on 1-0-1 (2) 3- Poa3- (b) F.C on P 0-1 Clo0 5- 6 4IC4]= 0 FC em 0gn (bCI03 Ce D O a 2-10 1-1 0 a F.C em 6-1-- b6- 6-1 [23 2 FCem 0ns = 6-4-L[43 6-6 = O 1 (1 11 C1 O 1 1/ CH3 CN H3CC N t-0-1 C8]=0